AN UPDATE ON THE VISIBLE REGION SPECTRUM OF YOH

K. ATHANASSENAS, C. T. KINGSTON, A. J. MERER, \underline{J. R. D. PEERS}, Chemistry Department, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada V6T 1Z1; S. J. RIXON, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada V6T 1Z1; D. A. GILLETT, Lambda Physik, Inc., 3201 West Commercial Blvd., Fort Lauderdale, FL 33309..

Further analysis of the visible region electronic band systems of YOH and YOD has been performed. Three electronic states, X̃^{1}Σ^{+}, B̃^{1}Π, and C̃^{1}Σ^{+}, have been found. Wavelength-resolved fluorescence spectra have given a large amount of information about the vibrational frequencies of the X̃^{1}Σ^{+} state. Exciting from the v′′_{2} =1 level of the ground state, we have been able to determine the separation of the l = 2 and 0 levels. The geometric structure of the ground state is found to be r(Y-O)= 1.94840(58) Å and r(O-H)=0.9215(53) Å. Due to the perturbed nature of the B̃^{1}Π, and C̃^{1}Σ^{+} states, an accurate determination of their geometries has not been possible. We have tentatively identified several members of the ν_{3} progressions in these states.